Grade 12Physical Chemistry

The radius of a divalent cation M2+ ion is 94 pm and of divalent anion X2-ion is 146 pm thus MX has

Yash Raj
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago

The cations and anions of an ionic precious stone are thought to be in contact with each other and henceforth the entirety of their radii will be equivalent to the bury atomic separation between them. 
r(C+) + r(A-) = d (C+-A-) 
where 
r(C+) - range of cation, C+ r(A-) - span of anion, A-= 94 – 126 = 32pm